### v2

@@ Methods
where we go from Equation 20 to Equation 21 by recognizing that randomly picking the number of positions to mask and sampling a random permutation is equivalent to picking a random time and sampling the forward noising process for that time. Thus, we have shown that the order-agnostic autoregressive modeling loss is equivalent, up to a constant, to a masking discrete flow matching loss with a uniform noise schedule.
Decoding ordering in order-agnostic autoregressive models
While the mathematical equivalence between order-agnostic autoregressive (OA-AR) models and discrete flow matching is exact in theory, a practical subtlety arises when implementing OA-AR models with causal attention mechanisms. In practice, many OA-AR models, including ProteinMPNN [1], are implemented using standard autoregressive architectures with causal attention masks. When computing the probability $p_{\theta}(x_{1}^{i}|x_{t})$ for a masked position $i$ given a partially masked sequence $x_{t}$, the model must process the unmasked positions in $x_{t}$ in some specific order due to the causal nature of the attention mechanism. Formally, for a position $i\in M_{t}$ (the set of masked positions), the probability $p_{\theta}(x_{1}^{i}|x_{t})$ should theoretically be independent of how the unmasked positions in $U_{t}={j:x_{t}^{j}\neq\text{?}}$ are ordered. However, in practice, the causal attention mechanism means that:
$p_{\theta}(x_{1}^{i}|x_{t})=p_{\theta}(x_{1}^{i}|x_{t}^{\sigma(1)},x_{t}^{\sigma(2)},\ldots,x_{t}^{\sigma(|U_{t}|)})$ (23)
where $\sigma$ is some permutation of the unmasked positions in $U_{t}$, and the probability depends on this specific ordering $\sigma$. During training, the OA-AR objective encourages the model to produce consistent predictions regardless of the permutation $\sigma$. The expectation over all possible permutations in the training objective should, in principle, ensure that $p_{\theta}(x_{1}^{i}|x_{t})$ is invariant to the ordering of unmasked positions. However, in practice, this invariance may not be perfectly achieved due to imperfect optimization and finite-sample training. To address this subtlety and ensure consistent behavior during guided sampling, for the protein stability experiment that guides ProteinMPNN, we fix the decoding ordering for $x_{t}$ based on the residue positions of the unmasked tokens. Specifically, we sort the unmasked positions in $U_{t}$ in ascending order of their sequence positions:
$\sigma(j)=\text{sort}(U_{t})[j]$ (24)
where $\text{sort}(U_{t})$ returns the unmasked positions ordered by their indices in the original sequence. This approach provides a deterministic and reproducible ordering for any given masked state $x_{t}$, respects the natural left-to-right order of residues in the protein sequence, and requires no additional hyperparameters or complex ordering strategies. By using this fixed position-based ordering, we ensure that the conditional probabilities $p_{\theta}(x_{1}^{i}|x_{t})$ are well-defined and consistent across different sampling runs, enabling reliable application of the guidance method to OA-AR models trained with causal attention.
Relationship to autoregressive models and discrete time diffusion models on discrete state-spaces
Autoregressive (AR) models represent another important class of generative models for discrete state-spaces that can be connected to the diffusion framework. AR models factorize the joint distribution sequentially as $p(x_{1})=\prod_{i=1}^{D}p(x_{1}^{i}|x_{1}^{1},x_{1}^{2},...,x_{1}^{i-1})$ , generating tokens one at a time in a fixed order. Austin et al. [19] showed that AR models can be interpreted as discrete space diffusion models with a particular form of noising process. For AR models, Yang and Klein [24] introduced FUDGE, a method that performs guidance by modulating each conditional distribution using a learned classifier evaluated on the partially completed sequence. FUDGE can be viewed as a special case of ProteinGuide when applied specifically to AR models.
While discrete time diffusion models [19] are trained with a different objective than their continuous-time counterparts, they can still benefit from our guidance methodology. The learned denoising model from a discrete-time training objective can be used to induce a rate matrix for sampling by assuming smooth interpolation between discrete time points. However, both theoretical analysis and empirical results [31, 33, 37, 38] suggest that models trained with continuous-time objectives often achieve superior performance in practice.
This relationship between different model classes highlights the generality of ProteinGuide as a framework that can accommodate various generative architectures, including AR models, discrete-time diffusion models in discrete state-spaces, and the continuous-time formulations we primarily focus on in this work.
Background on diffusion guidance
$\nabla_{x_{t}}\log p_{t}(x_{t}|y)=\nabla_{x_{t}}\log p_{t}(x_{t})+\nabla_{x_{t}}\log p_{t}(y|x_{t})$ (25)
$\nabla_{x_{t}}\log p^{(\gamma)}_{t}(x_{t}|y)=\nabla_{x_{t}}\log p_{t}(x_{t})+\gamma\nabla_{x_{t}}\log p_{t}(y|x_{t})$ (26)
$p(x_{t+dt}|x_{t},y)=\frac{p(y|x_{t+dt},x_{t})p(x_{t+dt}|x_{t})}{\sum\limits_{x^{\prime}_{t+dt}}p(y|x^{\prime}_{t+dt},x_{t})p(x^{\prime}_{t+dt}|x_{t})},$ (27)
where $x_{t+dt}$ and $x_{t}$ are separated by a infinitesimal time step $dt$, and $y$ is the desired property we wish to condition on. In a $D$-dimensional discrete state-space where each dimension has cardinality $S$, there are $S^{D}$ terms in the normalizing constant of Equation 27, rendering it generally intractable. Each term arises from the fact that, without constraints, any state can be reached from any other state. Intuitively, tractability can only be achieved if the number of terms is not exponential, thereby implying some constraints on the allowed state transitions. Next, we shall see how the continuous-time formulation of discrete state-space diffusion and flow models enables us to tractably compute the normalizing constant by imposing such constraints, without losing any model expressibility.
Guidance in discrete state-spaces
Here we provide a brief overview of the results presented in Nisonoff et al. [16]. To enable guidance for discrete state-space models, we turn to continuous-time diffusion models and flow models in discrete state-spaces. As described above, these models are based on continuous-time Markov chains (CTMCs) and therefore parameterize transition rates $R_{t}(x,\tilde{x})$ rather than score functions. These models also assume that the noising processes in each dimension are independent of one another (but not necessarily identical), which has the important consequence that the probability that two or more dimensions of the current state jump (transition) at exactly the same time is zero, which makes these models tractable [31, 37]. It turns out that the same assumption also makes guidance tractable for these models [16].
$R_{t}(x,\tilde{x}|y)=\frac{p(y|\tilde{x},t)}{p(y|x,t)}R_{t}(x,\tilde{x})$ (28)
$R_{t}^{(\gamma)}(x,\tilde{x}|y)=\left[\frac{p(y|\tilde{x},t)}{p(y|x,t)}\right]^{\gamma}R_{t}(x,\tilde{x})$ (29)
or derive predictor-free guidance discrete state-spaces:
$R_{t}^{(\gamma)}(x,\tilde{x}|y)=R_{t}(x,\tilde{x}|y)^{\gamma}R_{t}(x,\tilde{x})^{1-\gamma}$ (30)
$\log\frac{p_{\phi}(y|\tilde{x},t)}{p_{\phi}(y|x,t)}\approx(\tilde{x}-x)^{T}\nabla_{x}\log p_{\phi}(y|x,t)$ (31)
requiring only one forward and one backward pass of the predictor model instead of $D\times(S-1)+1$ forward passes, making estimation of the guide-adjusted rates $O(1)$ rather than $O(D\times S)$. With multiple particles, TAG can be further improved by acting as a proposal distribution in a Sequential Monte Carlo (SMC) algorithm similar to Wu et al. [48]. While there are other recently proposed approaches for guiding discrete diffusion models [49, 50, 51], ProteinGuide has the benefit of being both exact and general, applicable to both diffusion and flow-based models, making it a general and flexible framework for guiding a broad class of discrete state-space generative models.
Additional discussion of related work
Several recent works have proposed related but different techniques from Nisonoff et al. [16] for steering discrete state-space diffusion models. These approaches can be broadly categorized into guidance-based methods (which modify the sampling process at inference time) and fine-tuning approaches (which update the parameters of the pre-trained model).
Li et al. [49] proposed a derivative-free sampling technique for reward maximization in both continuous and discrete spaces using Sequential Monte Carlo (SMC). Their approach uses the unconditional model as the proposal distribution and reweights multiple particles with the reward function. Singhal et al. [51] built upon and generalized the methods in Wu et al. [48] and Li et al. [49] to enable more flexible choices over the sampling hyperparameters. Schiff et al. [50] presented a discrete-time analog of the continuous-time guidance method in Nisonoff et al. [16]. As noted in Schiff et al. [50], in discrete time, the normalizing constant in Equation 27 is still intractable, necessitating additional assumptions to achieve tractability, whereas Nisonoff et al. [16] showed that the continuous-time formulation of discrete diffusion models provides exactly the tractability necessary to perform exact guidance. In addition, the continuous-time formulation naturally enables guidance for discrete state-space flow matching models, whereas the discrete-time formation does not. Lee et al. [52] observed that when applying stronger guidance, sampling can be improved with SMC, aligning with challenges in low-temperature sampling noted in Du et al. [53] and Ingraham et al. [3]. For AR models, Zhao et al. [54] proposed an SMC-based steering method that learns intermediate twist functions via contrastive learning, similar to learning time-dependent predictors in diffusion guidance.
Recent works have also explored updating model parameters to incorporate conditioning rather than modifying the sampling process. Wang et al. [55] proposed to use the Gumbel-Softmax trick to make discrete diffusion trajectories differentiable and enable backpropagation of rewards through entire sampling paths. Concurrently, Rector-Brooks et al. [56] proposed to frame the steering problem as sampling from a Bayesian posterior where the pre-trained model serves as the prior and the reward acts as the likelihood. Both fine-tuning approaches established connections to Nisonoff et al. [16] while offering complementary perspectives on solving the steering problem.
Direct Preference Optimization (DPO [57]) and related work [11, 58] enable updating of pre-trained models, although up until recently, these methods have been applicable only to models with tractable likelihoods, thereby restricting them largely to autoregressive models. For example, Widatalla et al. [11] applied DPO to ESM-IF1, an autoregressive inverse folding model, to incorporate experimental stability data. In another example, Stocco et al. [59] applied DPO to an autoregressive model conditionally trained on enzyme classes to guide the generation towards various enzyme properties. Wallace et al. [60] extended DPO to diffusion models on continuous state-spaces by way of approximation, while Borso et al. [61] explored how to similarly do so for these models on discrete state-spaces. Importantly, the connections previously mentioned between MLMs, diffusion and flow matching models on discrete state-spaces, and order-agnostic autoregressive models, means that we can also generalize the aforementioned approximate DPO to these same model classes. Herein, we focus our study on plug-and-play guidance, rather than DPO-style approaches, although the latter could prove equally useful. Recent work by Yang et al. [62] compared various guidance-based and fine-tuning based approaches on several in-silico tasks for protein fitness optimization and found guidance-based approaches for discrete state-space diffusion models to be generally more effective than fine-tuning based approaches. In addition, they found that guidance-based approaches require less hyper-parameter tuning and have lower computational costs, making them more practically accessible to everyday users.
Reinforcement Learning from Human Feedback (RLHF) offers an alternative approach to model alignment through methods like Proximal Policy Optimization (PPO) [63]. PPO does not technically require the retrained model to have tractable likelihoods, but it requires an explicit reward model and on-policy training, making it often more complicated to train than more direct approaches like DPO [57]. In addition, the standard RLHF pipeline typically begins with more heuristic methods such as continuing to train an existing model with new data (i.e., supervised fine-tuning), for which the statistical outcome is not well-defined, and for which important information from the already-trained model may be forgotten. Blalock et al. [64] recently applied PPO to align protein language models e.g., ESM2 with experimental measurements. Notably, this approach does not directly sample from generative models in the conventional sense; instead, it employs a position-wise scoring strategy where models evaluate mutation probabilities at individual positions sequentially, followed by an iterative selection process that combines high-scoring mutations.
There is another line of work that model discrete data by embedding the data into continuous state-spaces, either on a probabilistic simplex [65] or on a continuous latent space [66, 67]. Some of the approaches proposed to apply guidance to diffusion models on the continuous state-space representations of discrete state-space objects in order to utilize the guidance approaches for continuous state-space diffusion. For example, Gruver et al. [68] propose to perform guidance in the hidden layers of the unconditional diffusion model by jointly training the unconditional denoising model and the classifier on one hidden layer of the denoising model. In another example, Lisanza et al. [69] propose to fine-tune RoseTTAFold with gaussian diffusion on the one-hot encoded sequence space in order to perform guidance in this continuous state-space. We note that approaches which perform guidance in continuous spaces lose the discrete structure of the data during generation, which can be important for settings where the discrete structures contain information that are useful for guidance [37].
Stability-guided inverse folding
We demonstrated how assay-labeled experimental data could be used to condition generative models for protein sequences, improving the success rates for such models in generating sequences with the desired function. To illustrate this use case, we guided ProteinMPNN with a large dataset of experimental folding stability measurements on mini-proteins [25] to generate sequences with improved stability. Stability is reported in real-valued units of $\Delta\Delta G$, which quantifies the difference in stability from the wild-type sequence. A protein is considered stable if it is at least as stable as the wild-type (i.e., $\Delta\Delta G\leq 0$). To generate sequences with improved stability, we followed the experimental setup from Nisonoff et al. [16]. Specifically, we used ProteinMPNN [1] as the inverse folding model, $p_{\theta}(\mathbf{s}\,|\,\mathbf{x})$, where $\mathbf{s}$ corresponds to a protein sequence and $\mathbf{x}$ is the desired backbone structure. As described above, the OA-AR nature of ProteinMPNN means we can treat the pre-trained model as a flow matching / diffusion model, $p_{\theta}(\mathbf{s}_{1}\,|\,\mathbf{s}_{t},\mathbf{x})$, under a masking noise process. Then we apply guidance with the stability predictor from Nisonoff et al. [16], $p_{\phi}(\Delta\Delta G\,|\,\mathbf{s},\mathbf{x})$, to obtain a stability-conditioned inverse folding model, $p_{\theta,\phi}(\mathbf{s}\,|\,\mathbf{x},\Delta\Delta G\leq 0)$.
We applied this set-up on the same eight randomly chosen proteins from Nisonoff et al. [16]. For each protein, we sampled 100 sequences using either ProteinMPNN or ProteinMPNN with guidance. For each generated sequence, we define success based on two criteria: (1) the generated sequence should fold into the desired structure, $\mathbf{x}$, and (2) the generated sequence should be at least as stable as the wild-type sequence, $\Delta\Delta G\leq 0$. We used AlphaFold 2 [70] to predict the structure of the generated sequences and compare the root mean square deviation (RMSD) between the predicted structure and the desired structure $\mathbf{x}$. We consider a sequence to achieve the desired fold if the RMSD is less than or equal to $2$Å following previous evaluation criteria [37]. To evaluate the folding stability of our sequences, we use the $\Delta\Delta G$ regression model $p(\Delta\Delta G\,|\,\mathbf{s},\mathbf{x})$. ProteinMPNN sampling was performed with three different temperatures: $1.0$, $0.1$, and $0.01$ and we report the results for the temperature that achieved the highest success rate. For guided generation, we used a single temperature of 0.1, Euler integration with a step size $\Delta t=0.005$, sampling stochasticity $\eta=20$ and guidance strength $\gamma=100$ for all proteins. On the whole, we observed that across all proteins, performing guidance with ProteinMPNN consistently improves its success rates (Fig. 2a). Details results are included in the Supplementary Information (Table A)
Enzyme class guided sequence generation
Overview
In this experiment, we demonstrated how guidance can generate sequences from a target enzyme classes under varying levels of sequence constraint. We trained an enzyme class predictor that classifies protein sequences into their Enzyme Commission (EC) numbers based on their sequence and used it to guide the protein language model ESM3 [4] during partial sequence redesign. EC numbers are hierarchical four-level codes describing different enzyme classes.
We use CLEAN [17] to evaluate the generated enzymes as an in silico benchmark [59]. This setup enabled us to explore the effectiveness of guidance with different sequence constraints, from complete sequence generation (no constraints) to heavily constrained redesign (50% sequence identity). We additionally present results for EC numbers of varying prevalence in the dataset, measuring generation success of the top-10 most common, 10 random, and 65 rare EC classes.
Enzyme class predictor
Our enzyme class predictor was a 3-layer multilayer perceptron (MLP) that took mean-pooled ESM3 sequence embeddings as input. The embeddings are of dimension 1536 and we used an MLP with hidden dimension ($2*1536=3072$).
The model was trained on the SwissProt dataset used by ProteInfer [71]. Unlike CLEAN [17], our model is trained on all SwissProt sequences, including non-enzymes. Sequences for each EC were randomly divided into training and test sets. During method development, we experimented with training classifiers initialized with the CLEAN weights, so the test set did not include any sequences from the CLEAN dataset. The test set contained sequences from ECs with at least 40 samples. At least 30 were used for training, and at most 10 went into the test set, as long as they weren’t in the CLEAN training data. Finally, 15% of the non-enzymes were included in the test set, and the rest were used as training data. The training data contained 315,833 samples and the testing data contained 36,354 samples. The noisy predictor’s dataset contained 10 randomly sampled noising times for each sequence.
The model was trained with a standard negative log-likelihood objective and an AdamW optimizer (learning rate of 0.001, weight decay coefficient of 0.01). We used a batch size of 15k and trained for two epochs on the full training data. The accuracy curves for each mask-level bin are plotted over the test set of the 100 most common level-4 EC numbers (Fig. 3).
Experimental setup
For each EC, we sampled 10 wild-type sequences (or as many as possible for rare ECs) from the test set and randomly masked positions according to the target masking level. Then, for each protein in our test set, we created partially masked sequences at 50-100% masked levels (50% is half the sequence fixed and 100% is generated from scratch).
In this experiment, we implemented exact classifier guidance for discrete diffusion as described in Nisonoff et al. [16] using the Gillespie sampling algorithm (Alg. 2, App. D from Peng et al. [39]). We used a guidance strength $\gamma=10$ on the classifier outputs and ESM3 sampling temperature of 0.7. We sampled all positions one at a time.
Evaluation metrics
We considered the generated sequences a success if CLEAN predicted they belonged to the target enzyme class and if they had a median pLDDT score of at least 0.8, according to the ESM3 structure prediction head.

### v4

@@ 1.3.2 Equivalence between MFMs and AO-ARMs
where we go from Equation (S27) to Equation (S28) by recognizing that randomly picking the number of positions to mask and sampling a random permutation is equivalent to picking a random time and sampling the forward noising process for that time. Thus, we have shown that the any-order autoregressive modeling loss is equivalent, up to a constant, to a masking discrete flow matching loss with a uniform noise schedule.
1.4 Simplified and faster sampling for flow matching and diffusion models with masked noise processes
In previous sections, we showed an equivalence between the loss functions used to train discrete flow-matching models (DFMs), discrete diffusion models (DDMs), any-order autoregressive models (AO-ARMs) and masked language models (MLMs), specifically when the DFMs and DDMs used a masked noise process, which we referred to as masked flow matching models (MFMs) and masked diffusion models (MDMs). In this section, we show a relationship between sampling algorithms used for these models.
1.4.1 Intuitive overview
We first provide an intuitive overview of the main technical result. We will show that instead of sampling from a CTMC—the required operation for sampling from a CTMC-based DFM—by way of integration in time using algorithms such as Euler, Gillespie, or $\tau$-Leaping (of which only Gillespie is exact), that one can instead use a much simpler algorithm to exactly sample from the CTMC. Specifically, one can use a simpler sampling algorithm that is normally used to sample from an AO-ARM. In particular, we will be able exactly sample from a CTMC underlying a DFM by independently sampling all transition times up-front—meaning before any state transitions have been sampled—and only then sampling all the state transitions. This lies in stark contrast to general CTMC sampling in which sampling time and state transitions must be interleaved. However, it is exactly how one would sample from an AO-ARM, where one first samples a ”decoding order“, and then autoregressively unmasks each position in turn according to this order, conditioning on all previously unmasked positions. Consequently, there is a constructive proof below for which AO-ARM to sample from so as to sample from the desired CTMC associated with a DFM. As a corollary, we also show that one only needs the ordering of the transition times to sample the state transitions.
In order to achieve correctness of this simplified sampling algorithm, we will need to make use of commonly used modeling choices made for CTMC-based DFMs, such as that these CTMCs follow an interpolation schedule as prescribed by the flow matching set-up. These choices enable us to derive a joint probability distribution for the CTMC over time and sequences that factorizes in the necessary way, namely into a term containing solely transition (holding) times, and another containing solely state transitions.
So far we have discussed only DFMs, but through other work that connects DFMs as a generalization of DDMs (see section below on the related work specific to sampling equivalances), these diffusion models inherit all the work of sampling from a DFM. We will discuss related work for DFMs below. Because we showed earlier that the training objective for MLMs is the same as for AO-ARM and MDMs, we can similarly appeal to our earlier reasoning of interpreting an MLM as either an AO-ARM or DFM, and correspondingly also sample from it in the same as described herein.
In addition to showing simplified and faster sampling for CTMC-based DFMs, we also, to the best of our knowledge, show for the first time how to use this new viewpoint to provide an alternative derivation for the guidance results for CTMC-based diffusion and flow matching models presented in Nisonoff et al. [82].
1.4.2 Related work on sampling algorithms
A number of existing works have discussed the topics herein, but these works are for the specific case of diffusion models, namely CTMC-based MDMs. We prove similar results for the case of flow matching models, namely CTMC-based MFMs. We also synthesize these results from prior work to give a coherent, constructive proof for the equivalence of paths sampled by MFMs and AO-ARM.
As mentioned above, it has been previously shown that sampling exactly from the CTMC of a MDM is equivalent to sampling unmasking times for each position upfront and then using this to define the decoding order to sample from an AO-ARM [86, 17, 49, 132, 83, 98, 94]. The discussions in Ou et al. [83] and Peng et al. [86] are most similar to this section in spirit, but again, analyze the equivalence for MDMs specifically. Furthermore, neither works give a unified proof of the equivalence in sampling which allows complete removal of the notion of time when sampling from an MFM, showing equivalence for the full path and not just marginal distributions. Appendix H of Campbell et al. [13] further shows a correspondence between DDMs and DFMs where DFMs define a large class of rate matrices for which DDMs correspond to a specific choice. We believe this holds in the general case, such that our results for MFMs should be reducible to the previously cited results for MDMs.
Previous works by Gat et al. [32], Campbell et al. [13] on MFMs have mentioned that the conditional distribution parameterized by the neural network is independent of time (see [94, 98, 86] for the analogous MDM results). This is a key insight that we use in this section as well. However, as used by prior work, we do not believe this insight suffices to show the AO-AR sampling method for MDMs samples the desired CTMC, nor does it directly extend the MDM proofs to MFMs. Herein, we provide a proof of the exact equivalence, which additionally gives a method to sample the transition times. Integrating these findings, Proposition 1 gives us a constructive proof to sample from the desired CTMC using either an MFM or an AO-ARM.
Our proofs hold for the common MFM formulation from Campbell et al. [13], Gat et al. [32], which we define again in the section below. Many of the results can readily be extended to other noise distributions, such as uniform substitutions, or to the use of stochasticity and corrector sampling. However, in those cases, a full reduction to AO-ARM sampling is not possible. Instead, the calculations yield an efficient Gillespie sampling algorithm, which we do not cover in detail.
1.4.3 Background and notation
In this section, we briefly review the relevant concepts for AO-ARM s, CTMCs, and flow-matching models. Although some of these were covered earlier, here we quickly remind the reader so as to also introduce notation convenient specifically to this section.
Sampling from an AO-ARM [112, 127, 106, 49] involves iteratively constructing a length-$D$ sample $X=(X^{1},\ldots,X^{D})$ starting from a fully masked sequence, in two steps. First, a permutation, $\sigma$, of positions, $1,2,\ldots,D$ is sampled uniformly at random. This permutation dictates the order of positions to unmask as the sample is constructed over $D$ iterations (one for each position), sometimes called the “decoding order”. Given the permutation ordering, $\sigma(1,2,\dots,D)$, the model progressively unmasks each position according to this ordering, one at a time, autoregressively. That is, at each step, the model samples from the conditional distribution, $X^{\sigma(i)}\sim p(X^{\sigma(i)}|x^{\sigma(<i)})$. These distributions are the core object learned when training an AO-ARM, typically by a neural network. We denote the set of progressively unmasked sequences generated during the autoregressive sampling as $\bm{X}=(X_{1},\ldots,X_{D})$, where bold indicates a ordered set of random variables, and subscript denotes the $i^{\text{th}}$ unmasking iteration. Note that both the dimension of $X$ is $D$, and that there are $D$ unmasking operations to generate it, which can be confusing. We sometimes refer to $\bm{X}$ as the “path”. We can also represent the path compactly as a tuple of the decoding order and the final sequence, denoted $\bm{X}=(\sigma,X)$.
Formally, we write the joint probability of the permutation ordering, $\sigma$, and the fully unmasked sequence, $X$ (together sometimes called the “decoding path”) as
$\mathbb{P}_{\mathrm{OA}}(\sigma,X)=U[S_{D}](\sigma)\prod_{i=1}^{D}p\left(X^{\sigma(i)}|x^{\sigma(<i)}\right),$ (S30)
where $S_{D}$ is the group of permutations over sequences of length $D$, $U[S_{D}](\sigma)$ indicates $\sigma$ is sampled uniformly at random from $S_{D}$, $X^{\sigma(i)}$ is the random variable for the value of position $\sigma(i)$, and $x^{\sigma(<i)}$ is the set of i1i-1 observed positions that are unmasked at the $i$-th decoding step.
Continuous time Markov chains (CTMCs) describe a sequence of Markovian transitions between discrete states as a continuous time process. For a detailed introduction, including path probability calculations like the ones done here, see Del Moral and Penev [23]. When a CTMC “jumps” from a state $x$ to state $\tilde{x}$ at time $\tau$, we say that the CTMC experienced a state transition from $x$ to $\tilde{x}$ at jump time $\tau$. Since the transition is instantaneous, we define the new state to be the state at time $\tau$: $X_{\tau}=\tilde{x}$. Accordingly, we say the previous state $x$ is the left limit of the CTMC at time $\tau$, meaning it is the value of the CTMC going right up to the jump time. We denote the left limit $X_{\tau_{i}}^{-}=\lim_{t\nearrow\tau_{i}}X_{t}=X_{\tau_{i-1}}$.
We will often omit the random variables when symbol for the observed value is the corresponding lower-case letter. For example, $\mathbb{P}_{\mathrm{MFM}}(x_{t})=\mathbb{P}_{\mathrm{MFM}}(X_{t}=x_{t})$ refers to the probability of observing the random variable $X_{t}$ taking on value $x_{t}$. Sometimes, when the variable for the observed value does not correspond to the random variable it is assigned to, we will write both explicitly. For example, in writing $\mathbb{P}_{\mathrm{MFM}}(x_{t}|X_{t}^{-}=x_{s})$ we refer to the probability that the masked flow matching model (MFM) transitions from $x_{s}$, the state it reached at a previous time $s$, to $x_{t}$ at time $t$. Recall that $X_{t}^{-}=x_{s}$ indicates that the state $x_{s}$ is the left limit $X_{t}^{-}$ of the state at $t$.
Finally, the rate matrix from $x$ to $\tilde{x}$ at time $t$ is denoted $R_{t}(x,\tilde{x}$). The rate matrix is generally time-dependent even when conditioned on the state. In flow-matching models, transition rates between states generally accelerate in time to ensure the process terminates by time $t=1$. Accordingly, the CDF of the jump time, is an integral of time-dependent rates and does not always have a closed-form expression. It is written as
$\mathbb{P}\left(T_{k}<t\mid T_{k-1}\right)=1-\exp\left(\int_{s=T_{k-1}}^{s=t}R_{s}\left(X_{s}^{-},X_{s}^{-}\right)\mathrm{d}s\right).$ (S31)
As one can see, this integrates the rate of the CTMC not undergoing a state transition at time $s$.
Discrete flow matching models (DFMs) were formulated by Campbell et al. [13], Gat et al. [32] to define a CTMC whose marginal distribution over states interpolates between a noise distribution $p_{\text{noise}}$ at time 0 and a data distribution $p_{\text{data}}$ at time $1$.
Formally, this requires that the rates learned by our DFMs are solutions to the Kolmogorov Forward Equation (KFE) that satisfy the interpolation $X_{t}=\kappa_{t}X_{1}+(1-\kappa_{t})X_{0}$ subject to $X_{1}\sim p_{\text{data}}$ and $X_{0}\sim p_{\text{noise}}$. The interpolation schedule $\kappa_{t}$ sets the probability of finding $X_{t}$ in the denoised state $X_{1}$ at each time $t$. The interpolation schedule can be freely defined by the user so long as $\kappa_{0}=0$ and it monotonically increases to $\kappa_{1}=1$. Note, for the MFMs analyzed in this section, this means we can interpret $\kappa_{t}$ as the CDF of the jump time for a position, $\kappa_{t}=\mathbb{P}(\tau<t)$. This, in turn, implies that the time derivative is the PDF of the jump time, $\dot{\kappa_{t}}=\mathbb{P}(\tau=t)$.
The CTMC underlying a DFM is used to sample transport paths to map samples from $p_{\text{noise}}$ to final samples from $p_{\text{data}}$, the distribution the DFM seeks to model. The transport is simulated by sampling a sequence $X_{0}\sim p_{\text{noise}}$ (all masked in the case of a masked flow-matching model, MFM) and numerically integrating the CTMC by sampling
$X_{t+\Delta t}\sim p_{t+dt|dt}(\cdot|x_{t})\approx\delta_{x_{t}}(\cdot)+R(x_{t},\cdot)\Delta t$
until we reach $X_{1}\sim p_{\text{data}}$.
The formulation of DFMs from Campbell et al. [13], Gat et al. [32] which was described above, makes three important assumptions: